OpenStudy (anonymous):

A 200.0mL flask of unknown gas was placed in a hot water bath. The temperature of the bath was found to be 99 degrees C. The pressure of the flask was 733mmHg. The flask with the unknown gas in it had a mass of 126.593g. The same flask weighed 125.623g when it was empty. Determine the mass f the unknown gas in the flask. Determine the Molar Mass of an unknown gas. (Assume that the Ideal Gas Law is obeyed and use R= 0.0821 L atm/mol K SHOW ALL WORK.)

OpenStudy (anonymous):

m(gas) = m(flask with a gas) - m(flask) m(gas) = 126.593 g - 125.623 g = 0.97 g P = 733mmHg = (733mmHg / (760 mmHg/atm) = 0.96atm d = m/V = 0.97g/0.2L = 4.85g/L R = 0.082 T = 99C + 273 = 372K by using PM=dRT 0.96 * M = 4.85*0.082*372 M = 154g/mol (look carefully, I may made a mistake..but I'm pretty sure I solved it right)